The following pension-related data pertain to Metro Recreation's noncontributory, defined benefit pension plan for 2018: ($ in 000s)

Jan. 1 Dec. 31

Projected benefit obligation $ 4,400 $ 4,680

Accumulated benefit obligation 3,730 3,980

Plan assets (fair value) 4,980   5,425

Interest (discount) rate, 6%

Expected return on plan assets, 10%

Prior service cost−AOCI (from Dec. 31, 2017, amendment) 870

Net loss−AOCI 518

Average remaining service life: 10 years

Gain due to changes in actuarial assumptions 44

Contributions to pension fund (end of year) 370

Pension benefits paid (end of year) 325

($ in 000s) PBO Plan Assets Prior Service Cost-AOCI 870 Cash Net Loss- AOCI 5181 Pension Expense 4,980 (4,400) (385) (264) NeRequired: Prepare a pension spreadsheet that shows the relationships among the various pension balances, shows the changes in those balances, and computes pension expense for 2018. (Enter credit amounts with a minus sign and debit amounts with a positive sign. Enter your answers in thousands (i.e. 200,000 should be entered as 200).)
Answer #1

($ in 000s)

PBO

Plan Assets

Prior Service Cost-AOCI

Net Loss AOCI

Pension expense

Cash

Net pension / (Liability) / Asset

Balance, Jan 1 2018

(4400)

4980

870

518

580

Service cost

(385)

385

(385)

Interest cost

(264)

264

(264)

Expected return on assets

498

(498)

498

Adjust for:

Loss on Assets

(98)

98

(98)

Amortization of:

Prior service cost

(87)

87

Net loss

(2)

2

Gain on PBO

44

(44)

44

Cash funding

370

(370)

370

Retiree benefits

325

(325)

Bal Dec 31 2018

(4680)

5425

783

570

240

745

Interest cost = 4400*6% = 264

Service cost = 4680-4400-264+44+325 = $385

Expected return on assets = 4980*10% = 498

Loss on assets = 498-400 = $98

Amortization of prior service cost = 870/10 = $87

Net loss = (518-498)/10 = 2